# African Data Ethics: A Discursive Framework for Black Decolonial Data Science

### v1

@@ Abstract.
Most artificial intelligence (AI) and other data-driven systems (DDS) are created by and for the benefit of global superpowers. The shift towards pluralism in global data ethics acknowledges the importance of including perspectives from the Global Majority to develop responsible data science (RDS) practices that mitigate systemic harms inherent to the current data science ecosystem. African practitioners, in particular, are disseminating progressive data ethics principles and best practices for identifying and navigating anti-blackness, colonialism, and data dispossession in the data science life cycle. However, their perspectives continue to be left at the periphery of global data ethics conversations. In an effort to center African voices, we present a framework for African data ethics informed by an interdisciplinary corpus of African scholarship. By conducting a thematic analysis of 47 documents, our work leverages concepts of African philosophy to develop a framework with seven major principles: 1) decolonize & challenge internal power asymmetry, 2) center all communities, 3) uphold universal good, 4) communalism in practice, 5) data self-determination, 6) invest in data institutions & infrastructures and 7) prioritize education & youth. We compare a subset of six particularist data ethics frameworks against ours and find similar coverage but diverging interpretations of shared values. We also discuss two case studies from the African data science community to demonstrate the framework as a tool for evaluating responsible data science decisions. Our framework highlights Africa as a pivotal site for challenging anti-blackness and algorithmic colonialism by promoting the practice of collectivism, self-determination, and cultural preservation in data science.
@@ 1. Introduction
Artificial intelligence (AI), machine learning (ML), and data science research and development are primarily conducted in economically powerful countries such as the United States, China, the United Kingdom, France, and Germany (Bengio et al., 2024). Consequently, data-driven systems (DDS) leveraging these methodologies often reflect the values of large corporations and Western ideals (Birhane et al., 2022). When ethical dilemmas arise within data science development, mitigation strategies are often limited by the perspectives of these actors who rarely experience the direct consequences of data harm (Mhlambi and Tiribelli, 2023; Cisse, 2018; Gwagwa et al., 2022).
@@
To balance the dominant Western perspective, pluralistic data ethics calls for perspectives from the Global Majority to better mitigate systemic concerns (Rifat et al., 2023; Carman and Rosman, 2023). Embracing pluralism in global data ethics acknowledges the limits of any single ethical perspective in guiding comprehensive responsible data science (RDS) practices. Unfortunately, African voices have routinely been left out of Western-centric data science ethics discussions (Eke et al., 2023c). This is a glaring omission—African data collaborators have an intimate experience with how colonialism, anti-blackness, and data dispossession operate in data science work (Abebe et al., 2021; Eke et al., 2023c). In addition, in the development of DDS, African people are often exploited as data workers, their resources are extracted for computing infrastructure, and even when deploying systems in their own communities, many components are managed and owned by external actors (Birhane, 2020).
@@
To address the underrepresentation of African perspectives in global data ethics, African data scientists have returned to African philosophies such as Ubuntu to articulate the current ethical dilemmas in data science work on the continent (Gwagwa et al., 2022). However, African data ethics texts are spread across a variety of publication venues and are rarely synthesized into a cohesive review or framework (Eke et al., 2022). For this reason, it can be difficult to grasp common topics, differing fields of thought, and how proposed principles translate to current practices of RDS in Africa.
In this work, we synthesize the current African data ethics discourse and contextualize data ethics theories through the analysis of African philosophies. We conduct a thematic analysis of 47 documents from African philosophy, information ethics, AI ethics, and human-computer interaction research to answer the following research question: What are the African values and ethical theories that can inform responsible data science (RDS) practices in Africa? As a result of our inquiry, we derived an ethical framework consisting of seven major principles rooted in the realities of data science work in Africa (Table 1). For example, one major principle we surfaced (decolonize & challenge internal power asymmetry) contends with the legacy of African colonialism and the current administration of neocolonialism to articulate the inefficacy of Western-centric data science in African contexts. Several principles also engage with a range of African communitarian theories to recontextualize common paradigms of social good, participatory design, and data ethics education.
@@
We place our framework in conversation with six other particularist data ethics frameworks to gauge its unique contributions to global efforts. While every framework is concerned with implementing technology for “common good”, we find African data ethics adds a pragmatic and nuanced perspective on how to uphold this ideal. To demonstrate the utility of our framework, we conclude with two case studies taken from ongoing African data science efforts. Overall, our work bridges rich African philosophical traditions with contemporary data science practices, developing a framework to guide responsible technology development within African contexts while contributing to global discussions on data ethics and justice.
@@ 2.1. Pluralistic Efforts in Global Data Ethics
Data ethics encompasses the normative frameworks and moral principles governing the collection, processing, storage, and deployment of data (Floridi and Taddeo, 2016). An ethical principle, in a data ethics context, represents a fundamental normative guide for action—such as respect for persons, beneficence, or justice—that transcends mere technical guidelines or professional best practices. What distinguishes genuine ethical principles from operational procedures is their foundation in moral philosophy and their universal applicability across contexts. Literature within this field has proposed new data ethics paradigms (Krijger et al., 2023; Verhulst, 2021; Burr and Leslie, 2023), critiqued current approaches to responsible data science (Taylor and Purtova, 2019; Leonelli, 2016; Gansky and McDonald, 2022), and experimented with algorithmic or other technical approaches to mitigate bias (Hort et al., 2024; Wang et al., 2023; Smith and Ricanek, 2020). While scholarly communities such as FAccT and AIES serve as platforms for engaging data ethics discourse, a significant portion of the scholarship produced in these venues has its roots in Western philosophy.
While Western philosophical traditions have dominated data ethics discourse through frameworks such as utilitarianism and deontological ethics (Morán-Reyes, 2022; Heavin and O’Connor, 2024), emerging scholarship increasingly recognizes the necessity of incorporating diverse ethical paradigms and epistemological frameworks from Global Majority perspectives. For example, Indigenous knowledge systems provide alternative conceptualizations of data sovereignty and stewardship that challenge Western individualistic notions of privacy and ownership (Carroll et al., 2023; Hudson et al., 2023; Walter et al., 2021; Lovett et al., 2019; Rainie et al., 2019). East Asian philosophical traditions, including Confucian ethics, contribute valuable insights regarding social harmony and collective responsibility in technological development, offering nuanced frameworks for balancing individual rights with communal interests in data-driven systems (D’Ambrosio, 2023; Ess, 2006; Yeh and Xu, 2010). Traditional African philosophy offers Ubuntu-based approaches that emphasize collective well-being and communal responsibility in data ethics (Reviglio and Alunge, 2020; Okyere-Manu, 2023).
@@ 2.2. African Philosophy: Foundation of African Data Ethics
African Philosophy is a sub-domain of philosophy meant to reclaim and generate philosophical theories from Sub-Saharan Africans (SSA) (Ramose, 2004b).111Throughout the rest of the document, Africa will refer to Sub-Saharan Africa. At the time of rapid decolonization, African philosophers pivoted from primarily engaging with Western philosophy to focus on uncovering the rich intellectual theories and practices of pre-colonial Africans (Wiredu, 2004a). In addition to examining the past, African philosophers work to cultivate an intellectual home for new philosophies that speak to the realities of modern African life (Ramose, 2004b; Kohnert, 2022).
Several African philosophical frameworks offer valuable insights for data ethics and responsible technology development. The Yoruba concept of ”iwa” (character/moral behavior) emphasizes the ethical implications of one’s actions on the collective community, suggesting approaches to data governance that prioritize communal benefit over individual gain (Oyeshile, 2021). Similarly, the Akan concept of ”onipa” (personhood) (Wingo, 2006; Owusu-Gyamfi, 2019) and the Zulu notion of Ubuntu, often translated as ”umuntu ngumuntu ngabantu” (a person is a person through other persons) (Reviglio and Alunge, 2020; Gwagwa et al., 2022), provides frameworks for understanding human dignity and agency in technological contexts, particularly relevant for issues of consent and data sovereignty. The Ethiopian philosophy of ”Medemer” (synergy/coming together) offers a model for collaborative data sharing and governance that balances individual autonomy with collective benefit (Assefa et al., 2024). Additionally, the Igbo concept of ”omenala” (customs/traditions) (Nwoye, 2011; Nwala, 1985) and the Swahili principle of ”ujamaa” (familyhood) (Nyerere, 1962) suggest that data governance should align with existing social structures and cultural practices rather than imposing external frameworks. Overall, these philosophies collectively lay the foundation for ethical frameworks that challenge Western individualistic approaches to data privacy, data sharing, and ownership while promoting interdependence and cultural alignment.
@@ 2.3. Current AI Ethics Discourse in Africa
A growing number of scholars are leveraging African philosophy to assess technological development and AI (Ewuoso, 2021; Capurro, 2008; Ruttkamp-Bloem, 2023; Gwagwa et al., 2022; Attoe et al., 2023). These works, along with foundational African philosophy, can inform practical approaches to data collection protocols that respect communal ownership and decision-making (Bühler et al., 2023), inspire algorithmic fairness metrics that incorporate African conceptions of justice and equity (Asiedu et al., 2024), and help shape privacy frameworks that balance individual rights with community interests (Jimoh, 2023). Additionally, African philosophy can be valuable in informing ethics review processes that consider local cultural contexts and values and help build data governance structures that reflect African leadership models and decision-making practices (Chinweuba, 2019). Our work aims to help bridge these issues and expand discourse on how African philosophies can contribute to advancing responsible data science practices within the continent.
@@ 2.3.1. Practitioner-led
Sábëlo Mhlambi’s pioneering work in African data ethics is grounded in the principles of Ubuntu (Mhlambi, 2020). He draws from the epistemic, ontological, and ethical theories of African philosophers like Mogobe Ramose (Ramose, 2004a) to critique prevailing understandings in AI ethics such as agency (Mhlambi and Tiribelli, 2023). Since Mhlambi’s Ubuntu data ethics contribution, a significant amount of African scholarship on data ethics highlights Ubuntu as an African philosophy that should be engaged with more in global data ethics work (Gwagwa et al., 2022; Segun, 2021; Langat et al., 2020; Kiemde and Kora, 2022; Goffi, 2023). This body of work makes a general appeal to African communitarian ethics to critique current ethical paradigms (Metz, 2021), report on data science work in Africa (Day and Rennie, 2023), and survey perceptions and concerns among African data science practitioners (Eke and Ogoh, 2022). There is also a plethora of African data ethics gray literature, information published outside of academic venues, published through African data organizations (often in partnership or funded by Western institutions) by way of blog posts (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023), reports (Sinha and Zulfa, 2023), and formalized briefs (Gwagwa, 2019). It must be noted that many contributions are made by Western, non-African data ethicists and disproportionately fail to include other African philosophies beyond Ubuntu.
@@ 3. Methods
Inspired by ongoing work, we have two aims in developing an ethical framework for African data ethics: 1) honoring the existing scholarship of African data ethicists by preserving diverging perspectives in the discourse, and 2) re-engaging with classic African philosophies to productively expand on African data ethics principles. To achieve the balance between a respectfully discursive and productively expansive framework, we combined methods from qualitative document analysis and literature reviews to develop our qualitative analysis protocol (Morgan, 2022; Birhane et al., 2022; Battle and Ottley, 2024).
@@ 3.1. Data Collection
The first author seeded the search by reading The African Philosophy Reader (Coetzee and Roux, 2004) due to prior exposure. This text informed subsequent keyword searches in established academic databases: Google Scholar, Web of Science, Scispace,222 Scispace is a language model search engine for literature reviews: https://typeset.io and the publication repositories of ACM and IEEE. The first author also searched their institutional library and online African philosophy libraries accessible through their institution. They used key phrases such as “African AI ethics”, “African philosophy” and “African data ethics” to identify relevant literature. Documents from database searches were excluded if African values and data science practices were not the primary topic. For example, Stahl et al. (2023) was excluded because its focus on North African data policy. Okolo et al. (2023) was included because the document primarily discussed the AI climate in SSA countries. All identified papers from the ACM Digital Library were ultimately excluded because African data values were often only discussed within a global overview of data ethics. In addition, documents were excluded if they were not a full document. Full documents were understood as non-archival and archival papers (no extended abstracts), reports, or book chapters. No range was set on the publishing year to permit the inclusion of foundational texts from African philosophy and African information ethics.
In parallel to keyword searches, we requested literature recommendations from other scholars in the field. In addition, we used reference and citation tracking to identify relevant documents that were missed in searches. By the end of our iterative data collection, 47 documents were collected. Details of inclusions/exclusions can be found in Appendix A.
@@ 3.2. Thematic Analysis
The first author reflexively coded the documents through a practice iteration followed by two rounds of coding to surface themes of African data ethics. First, the first author began a grounded reflection process by selecting six documents as a representative set of corpus topics (Coetzee, 2004b; Mhlambi, 2020; African Union, 2024; Ndjungu, 2020; Segun, 2021; Day and Rennie, 2023). From this selection, they highlighted and recorded meaningful excerpts from each document. Then they annotated a reflection about how the excerpt answers the research question. These reflections enabled the first author to focus the coding on what is considered an ethical principle. In subsequent coding iterations, we consider an African ethical principle to be a moral value, understanding, or standard prioritized by or derived from SSA communities. The first author then repeated the above process for each document in the full corpus, where they coded for statements that answer the research question. Then, they organized the resulting codes into themes through affinity diagramming (Scupin, 2008) using Miro. After the first full round of coding and principle clustering, several minor principles had conceptual gaps requiring the analysis of additional documents. For example, several documents mentioned centering women but lacked details in terms of responsible data science practices. To answer the open question, work from Pollicy333https://pollicy.org/ (an Afro-feminist technology collective based in Uganda) was collected to bolster the principle. In this iteration, excerpts the minor principles served as a codebook.

### v3

@@ Abstract.
The shift towards pluralism in global data ethics acknowledges the importance of including perspectives from the Global Majority to develop responsible data science practices that mitigate systemic harms in the current data science ecosystem. Sub-Saharan African (SSA) practitioners, in particular, are disseminating progressive data ethics principles and best practices for identifying and navigating anti-blackness and data colonialism. To center SSA voices in the global data ethics discourse, we present a framework for African data ethics informed by the thematic analysis of an interdisciplinary corpus of 50 documents. Our framework features six major principles: 1) Challenge Power Asymmetries, 2) Assert Data Self-Determination, 3) Invest in Local Data Institutions & Infrastructures, 4) Utilize Communalist Practices, 5) Center Communities on the Margins, and 6) Uphold Common Good. We compare our framework to seven particularist data ethics frameworks to find similar conceptual coverage but diverging interpretations of shared values. Finally, we discuss how African data ethics demonstrates the operational value of data ethics frameworks. Our framework highlights Sub-Saharan Africa as a pivotal site of responsible data science by promoting the practice of communalism, self-determination, and cultural preservation.
@@ 1. Introduction
Artificial intelligence (AI), machine learning (ML), and data science research and development are primarily conducted in economically powerful countries such as the United States, China, the United Kingdom, France, and Germany (Bengio et al., 2024). Consequently, data-driven technology (DDT) leveraging these methodologies often reflect the logics of large corporations and the hegemonic values of the countries they are based in (Birhane et al., 2022). When ethical dilemmas arise within data science development, mitigation strategies are often limited by the perspectives of these actors who rarely experience the direct consequences of data harm (Mhlambi and Tiribelli, 2023; Cisse, 2018; Gwagwa et al., 2022).
@@
To balance these dominant perspectives, pluralistic data ethics calls for perspectives from the Global Majority to better mitigate systemic concerns (Rifat et al., 2023; Carman and Rosman, 2023). Embracing pluralism in global data ethics acknowledges the limits of any single ethical perspective in guiding comprehensive responsible data science (RDS) practices. Unfortunately, Sub-Saharan African (SSA) 111Throughout the rest of the document, Africa or SSA will be usedACM FAcct 2025 Proceedings to refer to Sub-Saharan Africa. voices have routinely been left out of data science ethics discussions (Eke et al., 2023c). This is a glaring omission—African data collaborators have an intimate experience with how anti-blackness and data colonialism operate in data science work (Abebe et al., 2021; Eke et al., 2023c; Brevini et al., 2024). African people are often exploited as data workers, their resources are extracted for computing infrastructure, and even when deploying systems in their own communities, many processes are managed and owned by external actors (Birhane, 2020).
@@
To address the underrepresentation of African perspectives in global data ethics, African data scientists have returned to African philosophies such as Ubuntu to articulate the current ethical dilemmas in data science work on the continent (Gwagwa et al., 2022). However, African data ethics texts are spread across a variety of publication venues and are rarely synthesized into a cohesive review or framework (Eke et al., 2022). For this reason, it can be difficult to grasp common topics, differing fields of thought, and how proposed principles translate to current practices of RDS in Africa. In this work, we synthesize the current African data ethics discourse and contextualize data ethics theories through the application of African philosophies. We conduct a thematic analysis of 50 documents from African philosophy, information ethics, AI ethics, and human-computer interaction (HCI) research to answer the following research question: What are the African values and ethical theories that can inform responsible data science (RDS) practices in Africa? As a result of our inquiry, we derived an ethical framework consisting of six major principles rooted in the realities of data science work in Africa (Table 4).
Our framework surfaces novel understandings, theories, and practices for RDS in Africa that contend with the legacy of colonialism and the current administration of neocolonialism to articulate the inefficacy of current data science relations in African contexts. Several principles also engage with a range of African communitarian theories to recontextualize common paradigms of social good, participatory design, and comprehensive data ethics education. The framework also exhibits the practical solutions African data scientists offer to local governments, multinational technology companies, grassroots collectives, and individual data scientists engaged in the African data science ecosystem.
@@
We place our framework in conversation with other particularist data ethics frameworks to gauge its unique contributions to global efforts. While every framework is concerned with implementing technology for social good, we find African data ethics problematizes vague universalist language and calls for material redress for data harms. Overall, our work aims to bridge rich African philosophical traditions with contemporary data science practices, developing a framework that can guide responsible technology development within African contexts while contributing to global discussions on data ethics and justice.
@@ 2.1. Pluralistic Efforts in Global Data Ethics
Data ethics encompasses the normative frameworks and moral principles governing the collection, processing, storage, and deployment of data (Floridi and Taddeo, 2016). An ethical principle, in a data ethics context, represents a fundamental normative guide for action—such as respect for persons, beneficence, or justice—that transcends mere technical guidelines or professional best practices. What distinguishes genuine ethical principles from operational procedures are their foundation in moral philosophy and their universal applicability across contexts. Literature within this field has proposed new data ethics paradigms (Krijger et al., 2023; Verhulst, 2021; Burr and Leslie, 2023), critiqued current approaches to responsible data science (Taylor and Purtova, 2019; Leonelli, 2016; Gansky and McDonald, 2022), and experimented with algorithmic and other technical approaches to mitigate bias (Hort et al., 2024; Wang et al., 2023; Smith and Ricanek, 2020). While scholarly communities such as the ACM Conference oACM FAcct 2025 Proceedingsn Fairness, Accountability, and Transparency (FAccT) and the AAAI Artificial Intelligence, Ethics, and Society (AIES) serve as platforms for engaging discourse on AI and data ethics, a significant portion of the scholarship produced in these venues is primarily rooted in Western philosophy.
While Western philosophical traditions have dominated data ethics discourse through frameworks such as utilitarianism and deontological ethics (Morán-Reyes, 2022; Heavin and O’Connor, 2024), emerging scholarship increasingly recognizes the necessity of incorporating diverse ethical paradigms and epistemological frameworks from Global Majority perspectives. For example, Indigenous knowledge systems provide alternative conceptualizations of data sovereignty and stewardship that challenge Western individualistic notions of privacy and ownership (Carroll et al., 2023; Hudson et al., 2023; Walter et al., 2021; Lovett et al., 2019; Rainie et al., 2019). East Asian philosophical traditions, including Confucian ethics, contribute valuable insights regarding social harmony and collective responsibility in technological development, offering nuanced frameworks for balancing individual rights with communal interests in data-driven systems (D’Ambrosio, 2023; Ess, 2006; Yeh and Xu, 2010). Traditional African philosophy offers Ubuntu-based approaches that emphasize collective well-being and communal responsibility in data ethics (Reviglio and Alunge, 2020; Okyere-Manu, 2023). HCI and data ethics scholars have also applied knowledge from Black studies such as Black feminism (Reece et al., 2025), critical race theory (Coleman and Batiste, 2024), Afropessimism (Cunningham et al., 2023), prison abolitionism (Benjamin, 2019), and historical and contemporary social analysis (Gebru and Torres, 2024; Noble, 2018; BROWNE, 2015; Monroe-White, 2021) to articulate how data technology perpetuates anti-black racism in the United States.
@@ 2.2. African Philosophy: Foundation of African Data Ethics
African Philosophy is a sub-domain of philosophy meant to reclaim and generate philosophical theories from Africans (Ramose, 2004b). Around the time of widespread decolonization, African philosophers pivoted away from classic Western philosophy to uncover the rich intellectual theories and practices of pre-colonial Africans (Wiredu, 2004a). Beyond examining the past, African philosophers have also worked to cultivate an intellectual home for Africans to generate new philosophies that speak to the realities of African life in modernity (Ramose, 2004b; Kohnert, 2022).
Several African philosophical frameworks offer valuable insights for data ethics and responsible technology development. The Yoruba concept of “iwa” (character/moral behavior) emphasizes the ethical implications of one’s actions on the collective community, suggesting approaches to data governance that prioritize communal benefit over individual gain (Oyeshile, 2021). Similarly, the Akan concept of “onipa” (personhood) (Wingo, 2006; Owusu-Gyamfi, 2019) and the Zulu notion of Ubuntu, often translated as “umuntu ngumuntu ngabantu” (a person is a person through other persons) (Reviglio and Alunge, 2020; Gwagwa et al., 2022), provide frameworks for understanding human dignity and agency in technological contexts, particularly relevant for issues of consent and data sovereignty. The Ethiopian philosophy of “Medemer” (synergy/coming together) offers a model for collaborative data sharing and governance that balances individual autonomy with collective benefit (Assefa et al., 2024). Additionally, the Igbo concept of “omenala” (customs/traditions) (Nwoye, 2011; Nwala, 1985) and the Swahili principle of “ujamaa” (familyhood) (Nyerere, 1962) suggest that data governance should align with existing social structures and cultural practices rather than imposing external frameworks. Overall, these philosophies collectively lay the foundation for ethical frameworks that challenge Western individualistic approaches to data privacy, data sharing, and ownership while promoting interdependence and cultural alignment.
@@ 2.3. Current AI Ethics Discourse in Africa
A growing number of technology researchers are leveraging African philosophy to assess technological development and AI (Ewuoso, 2021; Capurro, 2008; Ruttkamp-Bloem, 2023; Gwagwa et al., 2022; Attoe et al., 2023). This work, along with foundational African philosophy, can inform practical approaches to data collection protocols that respect communal ownership and decision-making (Bühler et al., 2023), inspire algorithmic fairness metrics that incorporate African conceptions of justice and equity (Asiedu et al., 2024), and help shape privacy policies that balance individual rights with community interests (Jimoh, 2023). Additionally, African philosophy can be valuable in informing ethics review processes that consider local cultural contexts and values, and help build data governance structures that reflect African leadership models and decision-making practices (Chinweuba, 2019). Our work aims to help bridge these issues and expand discourse on how African philosophies can contribute to advancing responsible data science practices within the continent.
@@ 2.3.1. Practitioner-led
Sábëlo Mhlambi’s pioneering work on African data ethics, grounded in the principles of Ubuntu (Mhlambi, 2020), draws from the epistemic, ontological, and ethical theories of African philosophers like Mogobe Ramose (Ramose, 2004a) to critique prevailing AI paradigms (Mhlambi and Tiribelli, 2023). Since Mhlambi’s Ubuntu data ethics contribution, a significant amount of African scholarship on data ethics has highlighted Ubuntu as an African philosophy that should be engaged with more in global data ethics work (Gwagwa et al., 2022; Segun, 2021; Langat et al., 2020; Kiemde and Kora, 2022; Goffi, 2023). This body of work makes a general appeal to African communitarian ethics to critique current ethical paradigms (Metz, 2021), report on data science work in Africa (Day and Rennie, 2023), and survey perceptions and concerns among African data science practitioners (Eke and Ogoh, 2022). There is also a plethora of African data ethics gray literature published through African data organizations (often in partnership or funded by institutions in Europe or the United States) by way of blog posts (Shilongo, Kristophina and Peter, Kaulyaalalwa, 2023), reports (Sinha and Zulfa, 2023), and formalized briefs (Gwagwa, 2019). It must also be noted that many contributions made by non-African data ethicists disproportionately fail to engage with other African philosophies beyond Ubuntu.
@@ 3. Methods
Inspired by ongoing work, we have two aims in developing an ethical framework for African data ethics: 1) honoring the existing scholarship of African data ethicists by preserving distinct perspectives in the discourse, and 2) re-engaging with classic African philosophies to productively expand on African data ethics principles. To achieve the balance between a respectfully discursive and productively expansive framework, we combined methods from qualitative document analysis and literature reviews to develop our qualitative analysis protocol (Morgan, 2022; Birhane et al., 2022; Battle and Ottley, 2024).
@@ 3.1. Data Collection
The first author seeded the search by reading The African Philosophy Reader (Coetzee and Roux, 2004) due to prior exposure. This text informed subsequent keyword searches in established academic databases: Google Scholar, Web of Science, Scispace,222 Scispace is a language model search engine for literature reviews: https://typeset.io and the publication repositories of ACM and IEEE. The first author also searched their institutional library and online African philosophy libraries accessible through their institution. They used key phrases such as “African AI ethics”, “African philosophy” and “African data ethics” to identify relevant literature. Documents from database searches were excluded if African values and data science practices were not the primary topic. In addition, documents were excluded if they were not a full document. Full documents were understood as non-archival and archival papers (no extended abstracts), reports, or book chapters. No range was set on the publishing year to permit the inclusion of foundational texts from African philosophy.
In parallel to keyword searches, we requested literature recommendations from other scholars in the field. In addition, we used reference and citation tracking to identify relevant documents missed in searches. By the end of our iterative data collection, 50 documents were collected. The details of document inclusions and exclusions can be found in Appendix A.
@@ 3.2. Thematic Analysis
The first author reflexively coded the documents through a practice iteration followed by two rounds of coding to surface themes of African data ethics. First, the first author began a grounded reflection process by selecting six documents as a representative set of corpus topics (Coetzee, 2004b; Mhlambi, 2020; African Union, 2024; Ndjungu, 2020; Segun, 2021; Day and Rennie, 2023). From this selection, she highlighted and recorded meaningful excerpts from each document. Then she noted how the excerpt answers the research question. These reflections enabled the first author to focus on coding implicit or explicit African ethical principles. In subsequent coding iterations, we consider an African ethical principle to be a moral value, understanding, or standard prioritized by or derived from SSA communities. The first author then repeated the above process for each document in the full corpus, in which she coded excerpts that discussed African ethical principles relevant to RDS. Then, she organized the resulting codes into themes through affinity diagramming (Scupin, 2008) using Miro.333https://miro.com/ After the first full round of coding and principle clustering, she identified several minor principles with conceptual gaps that required the analysis of additional documents.
3.3. Positionality Statement
Our research team’s converging individual interests have led us to develop and propose our framework for African data ethics. The second author has extensive experience in developing AI policy in Africa and research expertise in developing explainable AI paradigms for marginalized communities. Other authors identify with a variety of African diasporic communities and are dedicated to expanding the representation of Black data science communities in their work on climate change, computer science education and responsible data science. For the author who does not belong to the African Diaspora, their work in pluralistic alignment of AI necessitates learning about the values of all user communities especially those on the margins. Since our research team is affiliated with academic and independent institutions based in the United States and the majority are not directly involved in technology collectives or disciplines represented in our corpus, the lens of epistemic humility guided our framework construction.
The first author is a Black, Jamaican-American woman. She acknowledges that her recent ethnicity is not Sub-Saharan African, but is committed to engaging and contributing to African data ethics. Her positionality as United States-educated, English monolinguist and observer of African data ethics discourse limits generative efforts. Her analytical perspective may also perpetuate misconceptions of the myriad of African perspectives highlighted in the work. As such, she aimed to reflexively guide the study by seeking guidance from technologists, policymakers and thought leaders within African data science communities.